# Inverse Problem Sampling in Latent Space Using Sequential Monte Carlo

### v2

@@ Abstract
In image processing, solving inverse problems is the task of finding plausible reconstructions of an image that was corrupted by some (usually known) degradation operator. Commonly, this process is done using a generative image model that can guide the reconstruction towards solutions that appear natural. The success of diffusion models over the last few years has made them a leading candidate for this task. However, the sequential nature of diffusion models makes this conditional sampling process challenging. Furthermore, since diffusion models are often defined in the latent space of an autoencoder, the encoder-decoder transformations introduce additional difficulties. To address these challenges, we suggest a novel sampling method based on sequential Monte Carlo (SMC) in the latent space of diffusion models. We name our method LD-SMC. We define a generative model for the data using additional auxiliary observations and perform posterior inference with SMC sampling based on a backward diffusion process. Empirical evaluations on ImageNet and FFHQ show the benefits of LD-SMC over competing methods in various inverse problem tasks and especially in challenging inpainting tasks.
@@ 1 Introduction
Figure 1: LD-SMC solves inverse problem tasks in the latent space of autoencoders by utilizing auxiliary observations ${\mathbf{y}}_{1:T}$ initialized using a forward diffusion process. Then, sampling ${\mathbf{z}}_{0}$ from the posterior distribution is done based on the backward diffusion process using sequential Monte Carlo. In the figure, ${\mathcal{D}}$ and ${\mathcal{A}}$ denote the decoder and the corruption operator respectively.
@@
Recently, several studies proposed sequential Monte Carlo (SMC) (Doucet et al., 2001b; Del Moral et al., 2012) as an effective solution for this task based on pixel-space diffusion models (Cardoso et al., 2023; Trippe et al., 2023; Wu et al., 2024; Dou & Song, 2024). Specifically, Wu et al. (2024) applies the DPS approximation (Chung et al., 2023b) for $p({\mathbf{y}}|{\mathbf{x}}_{t})$ with $p({\mathbf{y}}|\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}])$ and uses SMC sampling to correct for it. Dou & Song (2024), on the other hand, connected ${\mathbf{x}}_{t}$ to ${\mathbf{y}}$ by introducing a sequence of latent variables ${\mathbf{y}}_{1:T}$ through a duplex forward diffusion process and sampling sequentially from $p({\mathbf{x}}_{t:T}|{\mathbf{y}}_{t:T})$. While this approach has shown great potential, it has two main limitations. First, it does not take into account future observations ${\mathbf{y}}_{0:t-1}$ in the sampling process. Second, and more crucially, it is limited to linear corruption models only. As such, it cannot be applied with nonlinear mappings ${\mathcal{A}}$, let alone common Latent Diffusion Models (LDMs) (Rombach et al., 2022) due to the nonlinearity of the decoder. This is a harsh restriction as many of the recent powerful and efficient models are LDMs (Esser et al., 2024).
@@
Both existing approaches have pros and cons. Using the $p({\mathbf{y}}|\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}])$ approximation can be helpful in capturing the large-scale semantics of the image, but it often lacks in capturing the small details. On the other hand, using auxiliary observations ${\mathbf{y}}_{1:T}$ can help capture finer details, but the duplex forward diffusion process is not amenable to LDMs. Here, we propose a method that combines these two approaches and strives to achieve the best of both worlds. We define a generative model for the data based on a backward diffusion process, according to which an auxiliary observation ${\mathbf{y}}_{t}$ is generated directly from ${\mathbf{z}}_{t}$. Then, to apply posterior inference over the latent diffusion variables ${\mathbf{z}}_{0:T}$, we use SMC. To obtain a tractable sampling procedure, we derive an approximate target distributions and define a novel proposal distributions for the SMC sampling process tailored for diffusion models. Hence, we name our method Latent Diffusion Sequential Monte Carlo, or more concisely LD-SMC. An illustration of our approach is shown in Figure 1. We empirically validated LD-SMC on the ImageNet (Russakovsky et al., 2015) and FFHQ (Karras et al., 2019) datasets. We found that LD-SMC outperforms or is comparable to baseline methods on image deblurring and super-resolution tasks, and can significantly improve over baseline methods on inapainting tasks, especially on the more diverse ImageNet dataset.
@@
In this study, we make the following contributions: (1) we propose a novel method for combining auxiliary observations with latent space diffusion models; (2) we derive approximate target distributions for the SMC procedure and novel proposal distributions to perform approximate posterior sampling specifically tailored for diffusion models; (3) We theoretically show that LD-SMC is asymptotically accurate, namely that it can sample from $p_{\theta}(\mathbf{z}_{0}|\mathbf{y}_{0})$; (4) LD-SMC achieves significant improvements in perceptual quality in inpainting tasks, one of the most challenging inverse problem tasks. Our code is publicly available at https://github.com/ssi-research/LD-SMC
@@ 2 Background
[…]
@@ 3 Related Work
In (Song et al., 2021d) it was shown that to sample from the posterior distribution, $p({\mathbf{x}}_{0}|{\mathbf{y}})$, one can solve a stochastic differential equation based on the prior score, $\nabla_{{\mathbf{x}}_{t}}\log~{}p_{\theta}({\mathbf{x}}_{t})$, and the conditional score, $\nabla_{{\mathbf{x}}_{t}}\log~{}p_{\theta}({\mathbf{y}}|{\mathbf{x}}_{t})$. While the first term is easy to compute, the latter term requires integration over the full diffusion path from time $t$ to 00. A useful and easy-to-calculate approximation found in several studies is $p_{\theta}({\mathbf{y}}|{\mathbf{x}}_{t})\approx p_{\theta}({\mathbf{y}}|\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}])\approx p_{\theta}({\mathbf{y}}|\bar{{\mathbf{x}}}_{0}({\mathbf{x}}_{t}))$, which is readily available at each step (Chung et al., 2023b; Song et al., 2023; Wu et al., 2024). Specifically, Diffusion Posterior Sampling (DPS) (Chung et al., 2023b) uses this approximation for linear and nonlinear inverse problems with Gaussian and Poisson likelihood models. Other methods also utilize the pseudoinverse of the corruption operator ${\mathcal{A}}$ (Tirer & Giryes, 2018). $\Pi$GDM (Song et al., 2023) introduces a guidance scheme by matching the denoised output and the corrupted image ${\mathbf{y}}$, via a transformation of both. DDNM (Wang et al., 2023) suggested refining the contents of the null space of ${\mathcal{A}}$ during the backward diffusion process. A different approach proposed in (Mardani et al., 2024) is to approximate the posterior using a variational approach based on the score matching objective presented in (Song et al., 2021b). An additional category of inverse problem approaches that use diffusion models is designed with the objective of asymptotic exactness (Cardoso et al., 2023; Trippe et al., 2023; Wu et al., 2024; Dou & Song, 2024). These methods utilize SMC techniques, which have also been applied for unconditional sampling (Chen et al., 2025), to target exact sampling from the posterior distribution $p({\mathbf{x}}_{0}|{\mathbf{y}})$. SMC-Diff (Trippe et al., 2023) was designed mainly for motif-scaffolding. It uses the prior distribution as a proposal which is known to require a large number of particles for accurate estimation results (Särkkä, 2013), a severe limitation for latent space diffusion models due to expensive decoder evaluations. We experimented with this proposal in our initial attempts and witnessed that even $150$ particles were not enough to sample plausible reconstructions. MCGDiff (Cardoso et al., 2023) was designed for linear inverse problems only, which makes it unsuitable for inverse problems with latent-space diffusion models due to encoder-decoder involvement. TDS (Wu et al., 2024), a recent SMC-based method, solves general inverse problem tasks using the twisting technique. This method also uses the approximation of DPS, but by applying SMC sampling it can correct for it. FPS (Dou & Song, 2024), also a recent SMC-based method, uses auxiliary variables. FPS generates a sequence of observations ${\mathbf{y}}_{1:T}$ based on a duplex diffusion process, one process at the ${\mathbf{x}}$ space and the other process at the ${\mathbf{y}}$ space. Since this method is designed for linear inverse problems only, it permits tractable Bayesian inference. Our method combines the ideas of both TDS and FPS to obtain the best of both. Namely, we use the posterior mean approximation and ${\mathbf{y}}_{1:T}$ in our SMC sampling process. As we will show, this combination can be helpful in both understanding the general semantics of an image and capturing fine details. We note here that there are other methods that applied label augmentations similar to FPS, one example is (Abu-Hussein et al., 2022) which is applicable for linear inverse problems only. Hence, this method as well cannot be trivially combined with latent space diffusion models.
@@
1: Set ${\mathbf{z}}_{0}={\mathcal{E}}({\mathbf{x}}_{0})$, where ${\mathbf{x}}_{0}=\operatorname*{arg\,min}_{\mathbf{x}}||{\mathbf{y}}_{0}-{\mathcal{A}}({\mathbf{x}})||^{2}_{2}$
@@
Several inverse sampling methods were specifically tailored for latent diffusion models. PSLD (Rout et al., 2023) extend DPS (Chung et al., 2023b) by incorporating an additional gradient update step to guide the diffusion process to sample latent representations that maintain the integrity of the decoding-encoding transformation, ensuring it remains non-lossy. STSL (Rout et al., 2024) uses second-order information for reconstruction based on efficient approximations. Comparative analysis with STSL was not feasible due to the absence of publicly available code, making replication challenging. ReSample (Song et al., 2024), a contemporary method alongside PSLD, acts in two stages; it first applies hard data consistency to obtain latent variables that are consistent with the observed measurements and then employs a resampling scheme to map the sample back onto the data manifold. Concurrent to this study Nazemi et al. (2024) proposed a particle filtering approach. Their method builds on PSLD and DPS update in the proposal distribution. Similarly to TDS (Wu et al., 2024) the connection to the labels is only through ${\mathbf{z}}_{0}$ using the approximate mean estimator. Also, unlike LD-SMC it does not enjoy asymptotic guarantees. Recently, both DAPS (Zhang et al., 2024) and MGPS (Moufad et al., 2024) have been proposed to address inverse problems both in pixel and latent space. DAPS decouples latents in the sampling trajectory by first sampling from ${\mathbf{z}}_{0}|{\mathbf{z}}_{t+\Delta t},{\mathbf{y}}$ using MCMC techniques, for some $\Delta t>0$, and then sampling ${\mathbf{z}}_{t}$. MGPS also decomposes the sampling, but to an intermediate midpoint and uses variational inference for posterior sampling. Given the similarities of both methods, in our experiments we compare only to the former. SILO (Raphaeli et al., 2025), an additional recent study, first learns a NN that mimics an encoder operation on labels ${\mathbf{y}}$ which is then used to guide the sampling process. In our empirical evaluation, we compare to methods that apply sampling only.
@@ 4 Method
Given a corrupted image ${\mathbf{y}}_{0}$, the goal is to sample ${\mathbf{z}}_{0}\sim p_{\theta}({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$ using a pre-trained latent diffusion model as prior. Then, we can transform this sample into an image by applying a pre-trained decoder ${\mathcal{D}}$, i.e. ${\mathbf{x}}_{0}\coloneqq{\mathcal{D}}({\mathbf{z}}_{0})$. Here, we stack all the parameters of the networks (diffusion and encoder-decoder) under $\theta$. In what follows, we first define a generative model for the data and then apply Bayesian inference on all latent variables using blocked Gibbs sampling and SMC. Specifically, we use the diffusion process to augment the model with additional auxiliary observations and apply posterior inference using a sequential Monte Carlo (SMC) approach based on the backward diffusion process. The corresponding graphical model can be seen in Figure 2.
@@ 4.1 The Generative Model
$2.~{}{\mathbf{z}}_{t-1}|{\mathbf{z}}_{t}\sim p({\mathbf{z}}_{t-1}|{\mathbf{z}}_{t})\quad\forall t\in\{1,...,T\},$
$3.~{}{\mathbf{y}}_{t}|{\mathbf{z}}_{t}\sim\mathcal{N}({\mathcal{A}}(\underbrace{{\mathcal{D}}({\mathbf{z}}_{t})}_{{\mathbf{x}}_{t}}),\tau^{2}{\mathbf{I}})\quad\forall t\in\{0,...,{T}\}.$
Where, $p_{\theta}({\mathbf{y}}_{t}|{\mathbf{z}}_{t})$ is defined by the corruption model, $p({\mathbf{z}}_{t-1}|{\mathbf{z}}_{t})=\mathbb{E}_{p({\mathbf{z}}_{0}|{\mathbf{z}}_{t})}[p({\mathbf{z}}_{t-1}|{\mathbf{z}}_{t},{\mathbf{z}}_{0})]$ is a backwards generative process that corresponds to a non-Markovian forward diffusion process with
@@ 4.2.1 Initial Guess for ${\mathbf{z}}_{0}$
${\mathbf{x}}_{0}$ $=\operatorname*{arg\,max}_{\mathbf{x}}log~{}p_{\theta}({\mathbf{y}}_{0}|{\mathbf{x}})=\operatorname*{arg\,min}_{\mathbf{x}}||{\mathbf{y}}_{0}-{\mathcal{A}}({\mathbf{x}})||^{2}_{2},$ (3)
@@ 4.2.2 Posterior Sampling
$\begin{cases}p_{\theta}({\mathbf{y}}_{T}|{\mathbf{z}}_{T})\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{T})p({\mathbf{z}}_{T}),\qquad\qquad\qquad\qquad t=T,\\ \frac{p_{\theta}({\mathbf{y}}_{t}|{\mathbf{z}}_{t})\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{t})}{\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{t+1})}p_{\theta}({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})p_{\theta}({\mathbf{z}}_{t+1:T}|{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad~{}~{}0<t<T,\\ \frac{p_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{0})}{\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{1})}p_{\theta}({\mathbf{z}}_{0}|{\mathbf{z}}_{1})p_{\theta}({\mathbf{z}}_{1:T}|{\mathbf{y}}_{0:T}),\qquad\qquad\qquad t=0,\end{cases}$
where we define $\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{t})\coloneqq p_{\theta}({\mathbf{y}}_{0}|\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t}))=\mathcal{N}({\mathbf{y}}_{0}|{\mathcal{A}}({\mathcal{D}}(\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t}))),(1-\bar{\alpha}_{t}){\mathbf{I}}))$, with the variance taken to be the variance of ${\mathbf{z}}_{t}|{\mathbf{z}}_{0}$ according to the forward diffusion process. Importantly, as the final target distribution at $t=0$ matches the desired distribution $p_{\theta}({\mathbf{z}}_{0:T}|{\mathbf{y}}_{0:T})$ (see Eq. 10 in Appendix C), in the large compute limit samples from the correct target can be obtained despite all approximations.
@@
Inpainting (Box) Inpainting (Free-Form) Gaussian Deblurring Super-Resolution (8×8$\times$8 ×)
@@
The parameters $\gamma_{t}$ and $\lambda_{t}$ control the effect of the correction terms to the prior mean. In practice, it is challenging to control the trade-off between the two correction terms. Hence, we propose the following simple approach. During the sampling process, the first term only is used, that is, $\lambda_{t}=0$, and then starting from some predefined timestep $s$, the second term is used as well. The intuition here is that during the initial sampling steps, the quality of the labels ${\mathbf{y}}_{t}$ may not be good. Therefore, we rely on the first term through the posterior mean estimator to capture the general semantics of the image. However, in later sampling stages the quality of the labels increases (see Figure 7 in the appendix), and the latter correction term can help capture fine details in the image. This design choice also relates to the three-stage phenomenon in the diffusion sampling process witnessed in the literature (Yu et al., 2023). Setting $\lambda_{t}=0$ for all timesteps reduces the LD-SMC proposal update to that of TDS (Wu et al., 2024). Appendix D shows an instantiation of $\gamma_{t}$ and $\lambda_{t}$ used in this study.
@@
Connection to TDS. While the derivation is different, the SMC procedure in LD-SMC resembles somewhat that of TDS (Wu et al., 2024). The main difference between the two methods is the dependence on the auxiliary variables ${\mathbf{y}}_{1:T}$, which are only relevant for LD-SMC in both the proposal distributions and the weights. Empirically, we observed that this additional conditioning helped to better align the sampling with the corrupted image ${\mathbf{y}}_{0}$ compared to using the posterior mean approximation as in (Chung et al., 2023b) and (Wu et al., 2024).

### v3

@@ Abstract
In image processing, solving inverse problems is the task of finding plausible reconstructions of an image that was corrupted by some (usually known) degradation operator. Commonly, this process is done using a generative image model that can guide the reconstruction towards solutions that appear natural. The success of diffusion models over the last few years has made them a leading candidate for this task. However, the sequential nature of diffusion models makes this conditional sampling process challenging. Furthermore, since diffusion models are often defined in the latent space of an autoencoder, the encoder-decoder transformations introduce additional difficulties. To address these challenges, we suggest a novel sampling method based on sequential Monte Carlo (SMC) in the latent space of diffusion models. We name our method LD-SMC. We define a generative model for the data using additional auxiliary observations and perform posterior inference with SMC sampling based on a reverse diffusion process. Empirical evaluations on ImageNet and FFHQ show the benefits of LD-SMC over competing methods in various inverse problem tasks and especially in challenging inpainting tasks.
@@ 1 Introduction
Figure 1: LD-SMC solves inverse problem tasks in the latent space of autoencoders by utilizing auxiliary observations ${\mathbf{y}}_{1:T}$ initialized using the DDIM forward process. Then, sampling ${\mathbf{z}}_{0}$ from the posterior distribution is done based on the reverse diffusion process using sequential Monte Carlo. In the figure, $\mathcal{D}$} and $\mathcal{A}$} denote the decoder and the corruption operator respectively.
@@
Recently, several studies proposed sequential Monte Carlo (SMC) (Doucet et al., 2001b; Del Moral et al., 2012) as an effective solution for this task based on pixel-space diffusion models (Cardoso et al., 2023; Trippe et al., 2023; Wu et al., 2023; Dou & Song, 2024). Specifically, Wu et al. (2023) applies the DPS approximation (Chung et al., 2023b) for $p({\mathbf{y}}|{\mathbf{x}}_{t})$ with $p({\mathbf{y}}|\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}])$ and uses SMC sampling to correct for it. Dou & Song (2024), on the other hand, connected ${\mathbf{x}}_{t}$ to $\mathbf{y}$} by introducing a sequence of latent variables ${\mathbf{y}}_{1:T}$ through a duplex forward diffusion process and sampling sequentially from $p({\mathbf{x}}_{t:T}|{\mathbf{y}}_{t:T})$. While this approach has shown great potential, it has two main limitations. First, it does not take into account future observations ${\mathbf{y}}_{0:t-1}$ in the sampling process. Second, and more crucially, it is limited to linear corruption models only. As such, it cannot be applied with nonlinear mappings $\mathcal{A}$}, let alone common Latent Diffusion Models (LDMs) (Rombach et al., 2022) due to the nonlinearity of the decoder. This is a harsh restriction as many of the recent powerful and efficient models are LDMs (Esser et al., 2024).
@@
Both existing approaches have pros and cons. Using the $p({\mathbf{y}}|\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}])$ approximation can be helpful in capturing the large-scale semantics of the image, but it often lacks in capturing the small details. On the other hand, using auxiliary observations ${\mathbf{y}}_{1:T}$ can help capture finer details, but the duplex forward diffusion process is not amenable to LDMs. Here, we propose a method that combines these two approaches and strives to achieve the best of both worlds. We define a generative model for the data based on a reverse diffusion process, according to which an auxiliary observation ${\mathbf{y}}_{t}$ is generated directly from ${\mathbf{z}}_{t}$, the latent diffusion variable. Then, to apply posterior inference over the variables of all timesteps, ${\mathbf{z}}_{0:T}$, we use SMC. To obtain a tractable sampling procedure, we derive an approximate target distributions and define a novel proposal distributions for the SMC sampling process tailored for diffusion models. Hence, we name our method Latent Diffusion Sequential Monte Carlo, or more concisely LD-SMC. An illustration of our approach is shown in Figure 1. We empirically validated LD-SMC on the ImageNet (Russakovsky et al., 2015) and FFHQ (Karras et al., 2019) datasets. We found that LD-SMC outperforms or is comparable to baseline methods on image deblurring and super-resolution tasks, and can significantly improve over baseline methods on inapainting tasks, especially on the more diverse ImageNet dataset.
@@
In this study, we make the following contributions: (1) we propose a novel method for combining auxiliary observations with latent space diffusion models; (2) we derive approximate target distributions for the SMC procedure and novel proposal distributions to perform approximate posterior sampling specifically tailored for diffusion models; (3) We theoretically show that LD-SMC is asymptotically accurate, namely that it can sample from $p_{\theta}(\mathbf{z}_{0}|\mathbf{y}_{0})$; (4) LD-SMC achieves significant improvements in perceptual quality in inpainting tasks, one of the most challenging inverse problem tasks.
@@ 2 Background
${\mathbf{y}}_{T}$${\mathbf{y}}_{T-1}\cdots$${\mathbf{y}}_{1}$${\mathbf{y}}_{0}$${\mathbf{x}}_{T}$${\mathbf{x}}_{T-1}\cdots$${\mathbf{x}}_{1}$${\mathbf{x}}_{0}$${\mathbf{z}}_{T}$${\mathbf{z}}_{T-1}\cdots$${\mathbf{z}}_{1}$${\mathbf{z}}_{0}$
@@ 3 Related Work
[…]
@@
[…]
@@
Several inverse sampling methods were specifically tailored for latent diffusion models. PSLD (Rout et al., 2023) extend DPS (Chung et al., 2023b) by incorporating an additional gradient update step to guide the diffusion process to sample latent representations that maintain the integrity of the decoding-encoding transformation, ensuring it remains non-lossy. STSL (Rout et al., 2024) uses second-order information for reconstruction based on efficient approximations. Comparative analysis with STSL was not feasible due to the absence of publicly available code, making replication challenging. ReSample (Song et al., 2024), a contemporary method alongside PSLD, acts in two stages; it first applies hard data consistency to obtain latent variables that are consistent with the observed measurements and then employs a resampling scheme to map the sample back onto the data manifold. Concurrent to this study Nazemi et al. (2024) proposed a particle filtering approach. Their method builds on PSLD and DPS update in the proposal distribution. Similarly to TDS (Wu et al., 2023) the connection to the labels is only through ${\mathbf{z}}_{0}$ using the approximate mean estimator. Also, unlike LD-SMC it does not enjoy asymptotic guarantees. Recently, both DAPS (Zhang et al., 2024) and MGPS (Moufad et al., 2024) have been proposed to address inverse problems both in pixel and latent space. DAPS decouples latents in the sampling trajectory by first sampling from ${\mathbf{z}}_{0}|{\mathbf{z}}_{t+\Delta t},{\mathbf{y}}$ using MCMC techniques, for some $\Delta t>0$, and then sampling ${\mathbf{z}}_{t}$. MGPS also decomposes the sampling, but to an intermediate midpoint and uses variational inference for posterior sampling. Given the similarities of both methods, in our experiments we compare only to the former. SILO (Raphaeli et al., 2025), an additional recent study, first learns a NN that mimics an encoder operation on labels $\mathbf{y}$} which is then used to guide the sampling process. In our empirical evaluation, we compare to methods that apply sampling only.
@@ 4 Method
Given a corrupted image ${\mathbf{y}}_{0}$, the goal is to sample ${\mathbf{z}}_{0}\sim p_{\theta}({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$ using a pre-trained latent diffusion model as prior. Then, we can transform this sample into an image by applying a pre-trained decoder $\mathcal{D}$}, i.e. ${\mathbf{x}}_{0}\coloneqq{\mathcal{D}}({\mathbf{z}}_{0})$. Here, we stack all the parameters of the networks (diffusion and encoder-decoder) under $\theta$. In what follows, we first define a generative model for the data and then describe how to perform Bayesian inference on all latent variables using blocked Gibbs sampling. Specifically, we use the diffusion process and the corruption operator to augment the model with additional auxiliary observations. Then, inference is applied over the diffusion variables using SMC. The corresponding graphical model can be seen in Figure 2.
@@ 4.1 The Generative Model
$2.~{\mathbf{z}}_{t-1}|{\mathbf{z}}_{t}\sim p({\mathbf{z}}_{t-1}|{\mathbf{z}}_{t})\quad\forall t\in\{1,.,T\},$
$3.~{\mathbf{y}}_{t}|{\mathbf{z}}_{t}\sim\mathcal{N}({\mathcal{A}}(\underbrace{{\mathcal{D}}({\mathbf{z}}_{t})}_{{\mathbf{x}}_{t}}),\tau^{2}{\mathbf{I}})\quad\forall t\in\{0,.,{T}\}.$
Where, $p_{\theta}({\mathbf{y}}_{t}|{\mathbf{z}}_{t})$ is defined by the corruption model, $p({\mathbf{z}}_{t-1}|{\mathbf{z}}_{t})=\mathbb{E}_{p({\mathbf{z}}_{0}|{\mathbf{z}}_{t})}[p({\mathbf{z}}_{t-1}|{\mathbf{z}}_{t},{\mathbf{z}}_{0})]$ is a backwards generative process that corresponds to a non-Markovian forward process with
@@ 4.2.1 Initial Guess for ${\mathbf{z}}_{0}$
[…]
@@ 4.2.2 Posterior Sampling
[…]
where we define $\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{t})\coloneqq p_{\theta}({\mathbf{y}}_{0}|\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t}))=\mathcal{N}({\mathbf{y}}_{0}|{\mathcal{A}}({\mathcal{D}}(\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t}))),(1-\bar{\alpha}_{t}){\mathbf{I}}))$, with the variance taken to be the variance of ${\mathbf{z}}_{t}|{\mathbf{z}}_{0}$ according to the forward process. Importantly, as the final target distribution at $t=0$ matches the desired distribution $p_{\theta}({\mathbf{z}}_{0:T}|{\mathbf{y}}_{0:T})$ (see Eq. 10 in Appendix C), in the large compute limit samples from the correct target can be obtained despite all approximations.
@@
Inpainting (Box) Inpainting (Free-Form) Gaussian Deblurring Super-Resolution ($8\times$)
@@
The parameters $\gamma_{t}$ and $\lambda_{t}$ control the effect of the correction terms to the prior mean. In practice, it is challenging to control the trade-off between the two correction terms. Hence, we propose the following simple approach. During the sampling process, the first term only is used, that is, $\lambda_{t}=0$, and then starting from some predefined timestep $s$, the second term is used as well. The intuition here is that during the initial sampling steps, the quality of the labels ${\mathbf{y}}_{t}$ may not be good. Therefore, we rely on the first term through the posterior mean estimator to capture the general semantics of the image. However, in later sampling stages the quality of the labels increases (see Figure 7 in the appendix), and the latter correction term can help capture fine details in the image. This design choice also relates to the three-stage phenomenon in the diffusion sampling process witnessed in the literature (Yu et al., 2023). Setting $\lambda_{t}=0$ for all timesteps reduces the LD-SMC proposal update to that of TDS (Wu et al., 2023). Appendix D shows an instantiation of $\gamma_{t}$ and $\lambda_{t}$ used in this study.
@@
Connection to TDS. While the derivation is different, the SMC procedure in LD-SMC resembles somewhat that of TDS (Wu et al., 2023). The main difference between the two methods is the dependence on the auxiliary variables ${\mathbf{y}}_{1:T}$, which are only relevant for LD-SMC in both the proposal distributions and the weights. Empirically, we observed that this additional conditioning helped to better align the sampling with the corrupted image ${\mathbf{y}}_{0}$ compared to using the posterior mean approximation as in (Chung et al., 2023b) and (Wu et al., 2023).